OpenStudy (anonymous):

4 x (x + 2 ) =14 explain if the statement is correct or incorrect

OpenStudy (anonymous):

well you minus 2 from the four and you get 2x(x)=14 then you get 2 squared=14 and 2 squared is not 14 so it is incorrect.

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

thanks

OpenStudy (zzr0ck3r):

\[4x(x+2) = 14\\ 4x^2+8x = 14\\ 4x^2+8x-14=0\\ 4(x^2+2x)=14\\ 4(x+1)^2=14+4*1^2\\ 4(x+1)^2=18\\(x+1)^2=\frac{9}{2}\\x+1=\frac{\pm3}{\sqrt{2}}\\x=\frac{\pm3}{\sqrt{2}}-1\]

OpenStudy (zzr0ck3r):

@lowcard2 what you said maid no sense

OpenStudy (zzr0ck3r):

this is a quadratic equation, I am not quite sure what you were trying to do/

OpenStudy (anonymous):

everything that has been posted still makes no sense im confused

OpenStudy (anonymous):

I need to find what's wrong with the problem and correct it

OpenStudy (zzr0ck3r):

kitto are you sure this is exactly how they asked the question? you asked if a statement with a variable is true. this statement is only true for 2 x values, those are the ones I showed at the end... you could have gotten the same answers from \[ax^2+bx+c=0\\then\\x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (zzr0ck3r):

there is nothing wrong with the problem this is sort of like asking this... is this statement true 2+x = 4 well this is true when x = 2

OpenStudy (anonymous):

4 x (x + 2) = 14 4x + 2 =14 4x == 12 x =3

OpenStudy (zzr0ck3r):

you need to read about quadratic equations 4x(x+2) by the distribution propertiy of multiplication is 4x*x+2*4x 4x^2+8x so you have 4x^2+8x=14 we set equal to 0 and then solve 4x^2+8x-14=0 now we can use the formula I showed a = 4 b = 8 c = -14

OpenStudy (zzr0ck3r):

if this were true when x = 3 as you said then \[4*3(3+2) = 12(5) = 60\ne14\]